OpenStudy (anonymous):

If the potential difference between the plates of a capacitor is increased by 0.1%, the energy stored in the capacitor increases by exactly: A.) 20% B.) 22% C.) 40% D.) 44% Please show all the steps. The ans. should be D.

OpenStudy (anonymous):

the potential energy of a charged capacitor is \[W=CU^2/2\] where C is a capacitance and U is a potential difference. 1% of U is U/100, then 0.1% is 0.001U. So, if \[Unew=U+0.001U=1.001U\], then \[Wnew=C(1.001U)^2/2=(1.001)^2CU^2/2=(1.001)^2 W\]. It means tnat W is increased by \[(1.001)^2W-W=0.002001W\]. Hence it is increased by (0.002001)*100%=0.2001%. None of answers match!

OpenStudy (anonymous):

i have the same answer like kirill does. After doing reverse technique, i got that it's increasing of voltage is 20%. Maybe something need to be fix in the problem data if you want D as the answer... CMIIW

OpenStudy (anonymous):

the equaion for PE stored in a cpacitor is= 1/2 charge o one plate x te potentia differene